# CAST: Cross-Attentive Spatio-Temporal feature fusion for Deepfake detection

### v1

@@ Abstract
Deepfakes have emerged as a significant threat to digital media authenticity, increasing the need for advanced detection techniques that can identify subtle and time-dependent manipulations. CNNs are effective at capturing spatial artifacts and Transformers excel at modeling temporal inconsistencies. However, many existing CNN-Transformer models process spatial and temporal features independently. In particular, attention based methods often use independent attention mechanisms for spatial and temporal features and combine them using naive approaches like averaging, addition or concatenation, limiting the depth of spatio-temporal interaction. To address this challenge, we propose a unified CAST model that leverages cross-attention to effectively fuse spatial and temporal features in a more integrated manner. Our approach allows temporal features to dynamically attend to relevant spatial regions, enhancing the model’s ability to detect fine-grained, time-evolving artifacts such as flickering eyes or warped lips. This design enables more precise localization and deeper contextual understanding, leading to improved performance across diverse and challenging scenarios.We evaluate the performance of our model using the FaceForensics++, Celeb-DF, and DeepfakeDetection datasets in both intra and cross dataset settings to affirm the superiority of our approach. Our model achieves strong performance with an AUC of 99.49% and an accuracy of 97.57% in intra-dataset evaluations. In cross-dataset testing, it demonstrates impressive generalization by achieving a 93.31% AUC on the unseen DeepFakeDetection dataset. These results highlight the effectiveness of cross-attention-based feature fusion in enhancing the robustness of deepfake video detection.
@@ 1 Introduction
The rapid advancements in deepfake technologies fueled by advances in generative modeling and deep learning has made it increasingly easy to synthesize realistic facial videos that convincingly mimic real individuals. Such content, while promising for domains like film production, gaming, and virtual reality, poses severe threats to privacy, national security, and the integrity of digital media [1]. Consequently, deepfake detection has emerged as a crucial area of research to counter the misuse of these technologies [2]. Existing detection frameworks can be broadly categorized based on their reliance on spatial features, temporal dynamics, or a hybrid of both. Spatial feature-based approaches exploit frame-level artifacts, often introduced during generative synthesis. CNN-based models have been successful in identifying such localized inconsistencies, including texture irregularities and blending artifacts [3]. On the other hand, temporal models capture motion anomalies across frames, which are often introduced due to misalignment in expressions or inconsistencies in facial dynamics. Recurrent networks, 3D CNNs, and Transformer-based methods [4, 5] have shown potential in modeling such spatio-temporal cues. To improve detection robustness, hybrid spatio-temporal models have been proposed. These models typically leverage CNNs for extracting spatial features from each frame and Transformers or RNNs for temporal sequence modeling. In most cases, spatial and temporal branches are treated independently, with feature fusion performed using naive strategies such as concatenation, averaging, or late-stage score fusion [6, 7]. This limits the model’s capacity to model the intricate interplay between spatial cues and their temporal evolution. Recent works have attempted to address this limitation by integrating attention mechanisms to enhance spatial or temporal features. However, these mechanisms are often applied separately to the spatial and temporal domains, with limited cross-interaction. Among these methods, DFGaze [8] explored spatial-temporal gaze inconsistency, but avoided deep integration of spatial and temporal attention. Similarly, DDL [9] enhanced temporal modeling using an MSA module but avoided deep integration of spatial attention. In recent studies specifically, integrating multi-modal and cross-domain features has shown promise in enhancing generalization to unseen deepfake methods.
@@
In this paper, we propose a novel CAST hybrid model that leverages a cross-attention mechanism to dynamically fuse spatial and temporal features in a unified manner. Unlike traditional dual-stream approaches, our framework enables temporal tokens derived from transformer outputs to attend to spatial feature tokens extracted from CNNs, fostering rich inter-domain interactions. This dynamic fusion empowers the model to capture complex manipulations, such as flickering eyes or temporally inconsistent facial expressions, which may not be identifiable when modalities are processed in isolation. Our proposed model employs a CNN backbone to extract spatial features from individual frames. These features are then linearly projected and temporally aligned before being passed into a Transformer encoder, which captures temporal dependencies across the video sequence. A multi-head cross-attention module is introduced between the CNN-derived spatial embeddings and Transformer-encoded temporal tokens, enabling effective fusion of local spatial cues and global temporal context. Consequently, the architecture improves generalization across both intra- and cross-manipulation scenarios. Our key contributions can be summarized as follows:
@@
Author Year Category Feature Type Pre-processing Strength Weakness Qiu et al. [10] 2025 Hybrid Spatial and Frequency features Bi-directional and Spectral Attention with Feature Superposition Captures both local spatial and global frequency artifacts Lacks temporal modeling and relies on handcrafted frequency decomposition Wang et al. [11] 2022 Hybrid RGB and Frequency features Cross Modality Fusion (CMF) block Captures manipulation cues across spatial and frequency domains Lacks temporal modeling for video-based detection Siddiqui et al. [12] 2023 Hybrid Multi-scale + Attention CrossViT fusion Rich cross-modal representation High model complexity Yu et al. [13] 2023 Transformer Multi-view Spatiotemporal View-wise feature extraction + temporal alignment Strong cross-dataset generalization Requires high computational resources Wang et al. [14] 2023 Transformer Spatio-temporal Convolutional pooling Maintains coherence across frames Prone to overfitting Zhao et al. [15] 2021 Attention-based Face regions Multi-attention heads Localizes manipulated regions Sensitive to occlusions Face X-ray [3] 2020 Spatial Blending artifacts Region map extraction Fine-grained boundary detection Misses global inconsistencies Yoon et al. [16] 2024 Multimodal Facial, Audio, Text cues Triple-modality fusion Zero-shot generalization to unseen identities Higher computational cost Chen et al. [17] 2022 Self-supervised Adversarial features No labels needed Generalizes to unseen fakes May overfit synthetic noise
@@ 2.2 Deepfake Detection
Early works relied on spatial cues using CNNs like XceptionNet [20], which achieved high performance on FaceForensics++ [1]. Later hybrid models used CNN–LSTM or CNN–Transformer designs to capture spatio-temporal features. Temporal modeling using Vision Transformers has seen continued evolution. Recent work such as MSVT [13] proposed a transformer that captures multiple spatiotemporal views, allowing for robust generalization across varied manipulation methods.. With the rapid growth of deepfake generation techniques, a corresponding surge in deepfake detection methods has emerged to counter manipulated media, which is summarized in Table 1. These detection methods vary across spatial, temporal, spatio-temporal, multimodal, and attention-based approaches. Initially, CNN-based models were predominantly adopted for their capability to capture spatial inconsistencies within manipulated frames. Biswas et al. [21] utilized compact Xception framework with convolution block attention module to extract illumination invariant, chromaticity and depth features for the prediction of real and fake images. Rössler et al. [1] created the widely used FaceForensics++ dataset, providing a standardized benchmark for evaluating manipulated facial videos. Extending this, ForgeryNet [22] introduced a more diverse and large-scale dataset to facilitate robust training and cross-domain evaluation. These datasets helped reveal generalization issues in early detectors, especially when models trained on one manipulation type failed to detect others effectively. To improve generalizability, Li et al. [3] developed Face X-ray, a method capable of identifying blending artifacts between manipulated and authentic facial regions. This was extended by Zhao et al. [15] who used multi-attentional mechanisms to highlight discriminative facial regions. Siddiqui et al. [12] enhanced this approach using DenseNet and CrossViT to fuse multi-scale features and attention-based cues for robust classification. Wang et al. [11] proposed M2TR, which employs a Cross Modality Fusion (CMF) block to perform cross-attention based fusion between RGB and frequency features. Soudy et al. [23] introduced a hybrid deepfake detection approach that combines CNNs with a Vision Transformer, organizing the process into intuitive stages of preprocessing, targeted feature analysis, and final prediction through majority voting. Qiu et al. [10] proposed D2-Fusion, which combines bi-directional spatial attention and spectral frequency attention, followed by a feature superposition strategy to amplify artifact distinctions. Similarly, Wang et al. [14] adopted a convolutional pooling transformer to further enhance spatio-temporal coherence understanding in video forgeries. Multimodal detection has also gained popularity due to its ability to analyze non-visual inconsistencies. Yoon et al. [16] proposed TMI-former, a triple-modality deepfake detection framework that exploits facial, audio, and text modality interactions to detect forgeries in a zero-shot setting. Choi et al. [24] proposed a hybrid model combining a 3D ResNet with a StyleGRU module to capture both spatio-temporal features and latent style dynamics, improving deepfake detection generalization. Chen et al. [25] proposed a DeepFake detection framework based on Bi-Granularity Artifacts, capturing both intrinsic (e.g., upsampling) and extrinsic (e.g., blending) forgery traces. Their multi-task learning approach effectively exploits these artifacts to improve detection across diverse datasets. Tolosana et al. [26] also reviewed the broad landscape of face manipulation and fake detection, highlighting persistent challenges such as cross-domain generalization, real-time detection, and deepfake localization. Recent advancements have further enhanced the robustness of deepfake detection through hybrid and self-supervised approaches. Chen et al. [17] introduced a self-supervised learning method that leveraged adversarial examples for improved generalization ability of detection models and enabled better performance across varied manipulation types. Sun et al. [27] proposed a dual contrastive learning framework that aligns intra and inter-modality features using instance- and prototype-level losses, enhancing generalization under domain shifts. Zheng et al. [5] proposed a temporal coherence-based strategy to address the limitations of static frame analysis by enforcing consistency across sequential frames, thus improving the detection of subtle and temporally coherent forgeries.
@@ 3.1 Frame Selection and Preprocessing
where $\Delta$ denotes the frame sampling interval. This is necessary to ensure that frames are temporally spaced to capture diverse visual dynamics while maintaining computational feasibility.
@@
Each selected frame $I_{i}\in\mathbb{R}^{H\times W\times 3}$ is then passed through a face detection pipeline to localize facial regions of interest (ROI), which serve as the input to the downstream classification network.
3.1.1 Face Detection via MTCNN
For extracting facial regions under varying poses, scales, and occlusions, we utilize the Multi-Task Cascaded Convolutional Network (MTCNN), a three-stage cascaded deep learning architecture designed for face detection and facial landmark localization. MTCNN efficiently detects faces in images by leveraging a series of convolutional neural networks (CNNs), each focusing on progressively refining the quality of face localization and landmark detection. It consists of the Proposal Network (P-Net), the Refinement Network (R-Net), and the Output Network (O-Net). P-Net quickly proposes candidate face regions, R-Net eliminates false positives and refines bounding boxes. The overall MTCNN process is summarized by the equation:
$F_{i}$ $=\text{MTCNN}(I_{i})$ (3)
[…]
Here, each network in the MTCNN pipeline contributes to bounding box regression, classification, and non-maximum suppression (NMS), which progressively refine the candidate proposals and enhance face localization accuracy in diverse visual conditions.
@@ 3.4 Feature Fusion via Cross-Attention
$\mathcal{F}_{\text{fused}}=\text{LayerNorm}(\mathcal{Z}+\hat{\mathcal{Z}})$ (19)
@@ 4.1 Datasets
To comprehensively evaluate our proposed cross-attentive CNN-Transformer model for deepfake video detection, we conducted extensive experiments on three publicly available benchmark datasets: FaceForensics++ [1], DeepfakeDetection [31], and Celeb-DF (v2) [2]. These datasets encompass a wide range of manipulation techniques and visual qualities, enabling robust assessment across both intra- and cross-dataset scenarios.
@@
DeepfakeDetection (DFD): The DeepfakeDetection (DFD) dataset was released by Google and consists of over 3,000 high-resolution videos featuring both real and synthetically manipulated faces. These videos were created using multiple deepfake generation methods applied across various subjects, environments, and lighting conditions. DFD serves as a valuable benchmark for evaluating deepfake detection systems due to its realistic manipulations and diversity in terms of demographics, pose, and expressions. In our work, we utilized a curated subset of pristine and fake videos from DFD to conduct cross-dataset evaluations and multi-source training. The dataset’s diversity and naturalistic quality make it a robust benchmark for assessing generalization capability of detection models.

### v2

@@ Abstract
Deepfakes have emerged as a significant threat to digital media authenticity, increasing the need for advanced detection techniques that can identify subtle and time-dependent manipulations. CNNs are effective at capturing spatial artifacts and Transformers excel at modeling temporal inconsistencies. However, many existing CNN-Transformer models process spatial and temporal features independently. In particular, attention based methods often use independent attention mechanisms for spatial and temporal features and combine them using naive approaches like averaging, addition or concatenation, limiting the depth of spatio-temporal interaction. To address this challenge, we propose a unified CAST model that leverages cross-attention to effectively fuse spatial and temporal features in a more integrated manner. Our approach allows temporal features to dynamically attend to relevant spatial regions, enhancing the model’s ability to detect fine-grained, time-evolving artifacts such as flickering eyes or warped lips. This design enables more precise localization and deeper contextual understanding, leading to improved performance across diverse and challenging scenarios. We evaluate the performance of our model using the FaceForensics++, Celeb-DF, DeepfakeDetection and Deepfake Detection Challenge (DFDC) datasets in both intra and cross dataset settings to affirm the superiority of our approach. Our model achieves strong performance with an Area Under the Curve (AUC) of 99.49% and an accuracy of 97.57% in intra-dataset evaluations. In cross-dataset testing, the model achieves AUC scores of 93.31% and 81.25% on the unseen DeepFakeDetection and DFDC datasets, respectively. These results highlight the effectiveness of cross-attention-based feature fusion in enhancing the robustness of deepfake video detection.
@@ 1 Introduction
The rapid advancements in deepfake technologies fueled by advances in generative modeling and deep learning has made it increasingly easy to synthesize realistic facial videos that convincingly mimic real individuals. Such content, while promising for domains like film production, gaming, and virtual reality, poses severe threats to privacy, national security, and the integrity of digital media [33]. Consequently, deepfake detection has emerged as a crucial area of research to counter the misuse of these technologies [24]. Existing detection frameworks can be broadly categorized based on their reliance on spatial features, temporal dynamics, or a hybrid of both. Spatial feature-based approaches exploit frame-level artifacts, often introduced during generative synthesis. CNN-based models have been successful in identifying such localized inconsistencies, including texture irregularities and blending artifacts [22]. Recent studies in face-spoof detection [1, 2] demonstrate that spatial cue learning and deep feature representations can also distinguish authentic and manipulated facial content. On the other hand, temporal models capture motion anomalies across frames, which are often introduced due to misalignment in expressions or inconsistencies in facial dynamics. Recurrent networks, 3D CNNs, and Transformer-based methods [53, 52] have shown potential in modeling such spatio-temporal cues. To improve detection robustness, hybrid spatio-temporal models have been proposed. These models typically leverage CNNs for extracting spatial features from each frame and Transformers or RNNs for temporal sequence modeling. In most cases, spatial and temporal branches are treated independently, with feature fusion performed using naive strategies such as concatenation, averaging, or late-stage score fusion [25, 6]. This limits the model’s capacity to model the intricate interplay between spatial cues and their temporal evolution. Recent works have attempted to address this limitation by integrating attention mechanisms to enhance spatial or temporal features. However, these mechanisms are often applied separately to the spatial and temporal domains with limited cross-interaction. Among these methods, DFGaze [28] explored spatial-temporal gaze inconsistency, but avoided deep integration of spatial and temporal attention. Similarly, DDL [46] enhanced temporal modeling using an MSA module but avoided deep integration of spatial attention. In recent studies specifically, integrating multi-modal and cross-domain features has shown promise in enhancing generalization to unseen deepfake methods.
While several CNN– and Transformer–based architectures have been explored for deepfake detection, most existing methods still struggle to effectively combine spatial and temporal information in a unified manner. CNNs primarily focus on spatial artifacts such as blending boundaries or texture irregularities, yet they fail to capture frame-level temporal inconsistencies. Conversely, pure Transformer models require extensive data and computation to jointly learn spatial and temporal relations, often leading to limited generalization across unseen manipulations. The core problem this work addresses is the lack of a robust mechanism that can seamlessly fuse spatial and temporal representations while maintaining generalization and resilience to common real-world manipulations. To overcome these limitations, we propose a cross-attention-based spatial–temporal fusion framework (CAST) that enables the temporal stream to selectively attend to spatial cues extracted by the CNN, resulting in more discriminative and robust feature representations for deepfake detection.
@@
Unlike traditional dual-stream approaches, our framework enables temporal tokens derived from transformer outputs to attend to spatial feature tokens extracted from CNNs, fostering rich inter-domain interactions. This dynamic fusion empowers the model to capture complex manipulations, such as flickering eyes or temporally inconsistent facial expressions, which may not be identifiable when modalities are processed in isolation. Our proposed model employs a CNN backbone to extract spatial features from individual frames. These features are then linearly projected and temporally aligned before being passed into a Transformer encoder, which captures temporal dependencies across the video sequence. A multi-head cross-attention module is introduced between the CNN-derived spatial embeddings and Transformer-encoded temporal tokens, enabling effective fusion of local spatial cues and global temporal context. Consequently, the architecture improves generalization across both intra- and cross-manipulation scenarios. Our key contributions can be summarized as follows:
@@
Author Year Category Feature Type Pre-processing Strength Weakness Qiu et al. [31] 2025 Hybrid Spatial and Frequency features Bi-directional and Spectral Attention with Feature Superposition Captures both local spatial and global frequency artifacts Lacks temporal modeling and relies on handcrafted frequency decomposition Wang et al. [42] 2022 Hybrid RGB and Frequency features Cross Modality Fusion (CMF) block Captures manipulation cues across spatial and frequency domains Lacks temporal modeling for video-based detection Siddiqui et al. [35] 2023 Hybrid Multi-scale + Attention CrossViT fusion Rich cross-modal representation High model complexity Yu et al. [48] 2023 Transformer Multi-view Spatiotemporal View-wise feature extraction + temporal alignment Strong cross-dataset generalization Requires high computational resources Wang et al. [43] 2023 Transformer Spatio-temporal Convolutional pooling Maintains coherence across frames Prone to overfitting Zhao et al. [51] 2021 Attention-based Face regions Multi-attention heads Localizes manipulated regions Sensitive to occlusions Face X-ray [22] 2020 Spatial Blending artifacts Region map extraction Fine-grained boundary detection Misses global inconsistencies Yoon et al. [47] 2024 Multimodal Facial, Audio, Text cues Triple-modality fusion Zero-shot generalization to unseen identities Higher computational cost Chen et al. [5] 2022 Self-supervised Adversarial features No labels needed Generalizes to unseen fakes May overfit synthetic noise Hsu et al. [15] 2025 Diffusion / Inpainting Tiny partial fingerprint synthesis Keypoint-guided inpainting diffusion Generates 818K synthetic samples, 99.1% accuracy, strong augmentation Domain-specific to fingerprints, limited applicability to video modalities
@@ 2.1 Deepfake Generation
Early deepfake generation methods were primarily based on autoencoders and classical GAN models such as DeepFake-Autoencoder [20], enabling identity reconstruction from paired face images. With the introduction of advanced GAN architectures, particularly StyleGAN and StyleGAN2 [19, 17], synthesis quality dramatically improved by enabling high-resolution texture control and realistic facial structure manipulation. StyleGAN3 [18] later addressed temporal instability by eliminating aliasing artifacts, making generated video content visually more coherent.
Recently, diffusion-based generative models, such as DDPM [14] and Latent Diffusion [32], have demonstrated superior realism and semantic controllability compared to GANs through iterative denoising processes. These approaches have been increasingly adopted for face swapping and reenactment tasks under unconstrained conditions. Alongside diffusion models, transformer-based architectures have gained attention for deepfake generation, leveraging global self-attention to capture long-range dependencies more effectively than CNN-based designs. Vision Transformer-based generation frameworks [11] and masked autoencoder-based methods enable improved temporal coherence and semantic feature modeling across video frames.
In addition, multimodal deepfake generation frameworks combining face synthesis with speech and 3D modeling have emerged. Audio-driven reenactment models such as Wav2Lip [29] and SadTalker [50] generate realistic lip synchronization, while 3D NeRF-based reenactment approaches [27] produce high-fidelity facial dynamics and viewpoint control. Overall, the rapid evolution from autoencoders to GAN, diffusion, transformer, and multimodal frameworks has significantly increased the realism and complexity of manipulated media, reinforcing the necessity for robust detection models such as the proposed CAST framework.
@@ 2.2 Deepfake Detection
Early works relied on spatial cues using CNNs like XceptionNet [8], which achieved high performance on FaceForensics++ [33]. Later hybrid models used CNN–LSTM or CNN–Transformer designs to capture spatio-temporal features. Temporal modeling using Vision Transformers has seen continued evolution. Recent work such as MSVT [48] proposed a transformer that captures multiple spatiotemporal views, allowing for robust generalization across varied manipulation methods.. With the rapid growth of deepfake generation techniques, a corresponding surge in deepfake detection methods has emerged to counter manipulated media, which is summarized in Table 1. These detection methods vary across spatial, temporal, spatio-temporal, multimodal, and attention-based approaches. Biswas et al. [1] utilized compact Xception framework with convolution block attention module to extract illumination invariant, chromaticity and depth features for the prediction of real and fake images. Rössler et al. [33] created the widely used FaceForensics++ dataset, providing a standardized benchmark for evaluating manipulated facial videos. Extending this, ForgeryNet [13] introduced a more diverse and large-scale dataset to facilitate robust training and cross-domain evaluation. These datasets helped reveal generalization issues in early detectors, especially when models trained on one manipulation type failed to detect others effectively. To improve generalizability, Li et al. [22] developed Face X-ray, a method capable of identifying blending artifacts between manipulated and authentic facial regions. This was extended by Zhao et al. [51] who used multi-attentional mechanisms to highlight discriminative facial regions. Siddiqui et al. [35] enhanced this approach using DenseNet and CrossViT to fuse multi-scale features and attention-based cues for robust classification. Wang et al. [42] proposed M2TR, which employs a Cross Modality Fusion (CMF) block to perform cross-attention based fusion between RGB and frequency features. Soudy et al. [36] introduced a hybrid deepfake detection approach that combines CNNs with a Vision Transformer, organizing the process into intuitive stages of preprocessing, targeted feature analysis, and final prediction through majority voting. Qiu et al. [31] proposed D2-Fusion, which combines bi-directional spatial attention and spectral frequency attention, followed by a feature superposition strategy to amplify artifact distinctions. Similarly, Wang et al. [43] adopted a convolutional pooling transformer to further enhance spatio-temporal coherence understanding in video forgeries. Multimodal detection has also gained popularity due to its ability to analyze non-visual inconsistencies. Yoon et al. [47] proposed TMI-former, a triple-modality deepfake detection framework that exploits facial, audio, and text modality interactions to detect forgeries in a zero-shot setting. Choi et al. [7] proposed a hybrid model combining a 3D ResNet with a StyleGRU module to capture both spatio-temporal features and latent style dynamics, improving deepfake detection generalization. Chen et al. [4] proposed a DeepFake detection framework based on Bi-Granularity Artifacts, capturing both intrinsic (e.g., upsampling) and extrinsic (e.g., blending) forgery traces. Their multi-task learning approach effectively exploits these artifacts to improve detection across diverse datasets. Tolosana et al. [41] also reviewed the broad landscape of face manipulation and fake detection, highlighting persistent challenges such as cross-domain generalization, real-time detection, and deepfake localization. Recent advancements have further enhanced the robustness of deepfake detection through hybrid and self-supervised approaches. Chen et al. [5] introduced a self-supervised learning method that leveraged adversarial examples for improved generalization ability of detection models and enabled better performance across varied manipulation types. Sun et al. [38] proposed a dual contrastive learning framework that aligns intra and inter-modality features using instance- and prototype-level losses, enhancing generalization under domain shifts. Zheng et al. [52] proposed a temporal coherence-based strategy to address the limitations of static frame analysis by enforcing consistency across sequential frames, thus improving the detection of subtle and temporally coherent forgeries. Hsu et al. [15] proposed an Inpainting Diffusion Synthetic approach guided with feature keypoints for generating tiny partial fingerprint samples, producing over 818,077 synthetic images and achieving up to 99.1% matching accuracy. Their work demonstrates the significance of high-quality data augmentation for improving recognition robustness.
@@ 3.1 Frame Selection and Preprocessing
where $\Delta$ denotes the frame sampling interval. This ensures that frames are temporally spaced to capture diverse visual dynamics while maintaining computational feasibility.
@@
Each selected frame $I_{i}\in\mathbb{R}^{H\times W\times 3}$ is then processed through a face detection and alignment step using the Multi-task Cascaded Convolutional Network (MTCNN) [49], which ensures consistent facial region cropping and alignment across varying poses and lighting conditions. The detected face regions constitute the regions of interest (ROI) fed to the downstream classification network.
@@ 3.4 Feature Fusion via Cross-Attention
[…]
@@
In the proposed cross-attention fusion, temporal tokens are used as queries while spatial tokens operate as keys and values. This formulation reflects the functional relationship between temporal and spatial information in deepfake analysis. Temporal dynamics frequently expose inconsistencies such as motion discontinuities, lip-sync mismatch, or frame-level blending artifacts, which are strong indicators of manipulation. By treating temporal features as queries, the model effectively asks: “Which spatial regions are relevant given the current temporal context?” In contrast, spatial features represent a more stable repository of structural appearance cues, functioning as a spatial memory from which values are selectively retrieved based on learned relevance. The attention weights compute similarity between temporal queries and spatial keys, and the resulting value aggregation enhances spatial details precisely where temporal cues detect anomalies. Empirically, our directionality ablation study (Sec. 5.5.3, Table 6) demonstrates that reversing the Q/K roles leads to reduced cross-dataset AUC performance, supporting this asymmetric design choice.
@@ 4.1 Datasets
To comprehensively evaluate our proposed cross-attentive CNN-Transformer model for deepfake video detection, we conducted extensive experiments on three publicly available benchmark datasets: FaceForensics++ [33], DeepfakeDetection [12], Celeb-DF (v2) [24] and Deepfake Detection Challenge [10]. These datasets encompass a wide range of manipulation techniques and visual qualities, enabling robust assessment across both intra- and cross-dataset scenarios.
@@
DeepfakeDetection (DFD): The DeepfakeDetection (DFD) dataset was released by Google and consists of over 3,000 high-resolution videos featuring both real and synthetically manipulated faces. These videos were created using multiple deepfake generation methods applied across various subjects, environments, and lighting conditions. DFD serves as a valuable benchmark for evaluating deepfake detection systems due to its realistic manipulations and diversity in terms of demographics, pose, and expressions. In our work, we utilized a curated subset of pristine and fake videos from DFD to conduct cross-dataset evaluations. The dataset’s diversity and naturalistic quality make it a robust benchmark for assessing generalization capability of detection models.
@@
Deepfake Detection Challenge (DFDC): The Deepfake Detection Challenge (DFDC) dataset is a large-scale benchmark for deepfake video detection, comprising of 19,154 real videos and 100,000 manipulated videos generated from 3,426 compensated actors under controlled and unconstrained recording conditions. The manipulated videos were created using multiple deepfake generation pipelines, resulting in diverse visual artifacts, identity substitutions, and temporal inconsistencies. The dataset exhibits substantial variability in facial attributes, head poses, illumination conditions, and background scenes, making it particularly challenging for deepfake detection. It introduces significant distributional shifts and weak manipulation cues that closely resemble real-world forgery scenarios. In this work, DFDC is primarily used for cross-dataset evaluation to assess the generalization capability of the proposed model across diverse data distributions and manipulation patterns.